x/4-y/11=-9/44
x/6-y/13=-1/26

Answer by stanbon(75887)

x/4-y/11=-9/44
x/6-y/13=-1/26
----
Multiply thru the 1st equation by 44
Multiply thru the 2nd equation by 6*26
---------------------------
11x - 4y = -9
26x - 12y = -6
--------------------
Multiply thru the 1st equation by 3
33x -12y = -27
26x -12y = -6
------------------
Subtract and solve for "x":
7x = -21
x = -3
------------
Solve for "y":
11x - 4y = -9
11*-3 - 4y = -9
-33 -4y = -9
4y = -24
y = -6
===================
Solution: (-3,-6)
========================
Cheers,
Stan H.